OpenStudy (anonymous):

determine if the given sequence is an arithmetic sequence, if so find the common difference n,1,2,3,4 f(n),-3,-2,-1,0

OpenStudy (anonymous):

0--1=1 -1--2 = 1 -2--3=1 -3--4 =1 so there is a common difference of terms